Question 5 (1 point) Savea The revenue of 200 companies is plotted and found to follow a bell curve. The mean is $582.869 million with a standard deviation of $34.5316 million. Would it be unusual for a randomly selected company to have a revenue below $445.71 million? O 1) We do not have enough information to determine if the value is unusual 2) The value is borderline unusual. 3) The value is not unusual. 4) The value is unusual 5) It is impossible for this value to occur with this distribution of data
Answer #1

this will be your range of usual:

=(34.5316*2)+582.869 = 651.93 this is your highest value

=582.869-(32.5319*2) = 517.80 this is your lowest value

Usual range: 517.80 ≤ x ≤ 651.93

Since 445.71 is less than your lowest end, 517.80, then it is definitely unusual.

so the answer is , option 4

The value is unusual

The formal definition of unusual is a data value more than 2 standard deviations away from the mean in either the positive or negative direction